# Burst-tree structure and higher-order temporal correlations

### v1

@@ Abstract
Understanding characteristics of temporal correlations in time series is crucial for developing accurate models in natural and social sciences. The burst-tree decomposition method was recently introduced to reveal higher-order temporal correlations in time series in a form of an event sequence, in particular, the hierarchical structure of bursty trains of events for the entire range of timescales [Jo et al., Sci. Rep. 10, 12202 (2020)]. Such structure has been found to be simply characterized by the burst-merging kernel governing which bursts are merged together as the timescale for detecting bursts increases. In this work, we study the effects of kernels on the higher-order temporal correlations in terms of burst size distributions, memory coefficients for bursts, and the autocorrelation function. We employ several kernels, including the constant, additive, and product kernels as well as those inspired by the empirical results. We find that kernels with preferential mixing lead to the heavy-tailed burst size distributions, while kernels with assortative mixing lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the interevent time distribution. In addition, thanks to the analogy to the coagulation process, analytical solutions of burst size distributions for some kernels could be obtained.
@@ I Introduction
Recently, Jo et al. introduced a novel method called burst-tree decomposition to analyze temporal correlations in the event sequence [20]. This method represents an event sequence in terms of the burst tree, which is decomposed into a set of IETs and an ordinal burst tree. The ordinal burst tree exactly captures the hierarchical structure of correlations between IETs in a systematic way. It turns out that several time series data from diverse backgrounds with different IET distributions show the similar burst-tree structure in terms of heavy-tailed burst size distributions and positive correlations between consecutive burst sizes for a wide range of timescales. The latter was quantified in terms of a memory coefficient for bursts. The original definition of the memory coefficient has been proposed for IETs [17]. The burst-tree structure could be simply summarized by a burst-merging kernel governing which bursts are merged together as the timescale increases. The estimated burst-merging kernels from the data have revealed the preferential and assortative mixing structure of bursts. Then a model adopting a simple mathematical form of the burst-merging kernel based on the empirical analysis was studied to successfully generate temporal correlations in the original time series, providing a powerful tool for modeling and analyzing complex systems.
@@
The modeling approach using burst-merging kernels is analogous to the coagulation process by which unit particles are merged according to some selection rules or kernels, to become clusters [33, 34, 35, 36, 37, 38, 39]. Different approaches have been taken to derive cluster size distributions as well as power-law exponent values in the cases of power-law distributions of cluster sizes. In particular, Wattis employed a discrete deterministic mean-field approach to derive the analytical solution of the cluster size distribution for three solvable kernels, such as constant, additive, and product kernels [39]. Since unit particles and clusters in the coagulation process respectively correspond to individual events and bursts in the event sequence, the analytical solutions of cluster size distributions in the coagulation process can be translated into the burst size distributions of the event sequence generated using the same kernels. It implies that the analytical solutions of cluster size distributions in Ref. [39] can be used to confirm the burst size distributions obtained when the constant, additive, and product kernels are employed for the burst-merging kernels.
@@
Despite the importance of burst-merging kernels in time series analysis, their impact on higher-order temporal correlations remains poorly understood. To address this issue, we conduct a systematic comparison of various kernels, including the constant, additive, and product kernels as well as those inspired by empirical results. By analyzing the event sequences generated using these kernels and power-law IET distributions, we investigate burst size distributions, memory coefficients for bursts, and ACFs. We generically find that kernels with preferential mixing lead to the heavy-tailed burst size distributions, while kernels with assortative mixing lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the IET distribution. These findings provide new insights into the role of burst-merging kernels in shaping the statistical properties of time series, and set the stage for the remainder of this paper, which delves deeper into the implications of these results.
@@ II Burst-tree decomposition method
The burst-tree decomposition method [20] was recently introduced to systematically reveal temporal correlations in time series, in particular, the hierarchical structure of bursts for a wide range of timescales defining the bursts. A burst refers to a sequence of events that occur in rapid succession, followed by a period of relative inactivity [8]. To be precise, for a given timescale $\Delta t$, a burst is defined as a set of events where interevent times (IETs) between consecutive events in the burst are less than or equal to $\Delta t$, whereas the IETs between bursts are greater than $\Delta t$ [22]. It implies that two consecutive bursts detected using a timescale $\Delta t$ would be merged for a bigger timescale $\Delta t^{\prime}$ ($>\Delta t$) if $\Delta t^{\prime}$ is larger than the IET separating these two bursts. Such merging pattern is captured by the burst-tree decomposition method.
@@
Since the burst tree contains exactly the same information in the event sequence, one can recover the original event sequence $\mathcal{E}$ from the derived burst tree $\mathcal{T}$ too. This is done by performing an inorder traversal of the burst tree. If the $i$th visited internal node is indexed by $u$, we denote it by $u(i)$. Then the event timings are computed as $t_{i}=t_{i-1}+\hat{\tau}_{u(i)}$. Here $t_{0}$ can be taken from the time series data, or assumed to be 00 for simplicity. Thanks to the derivability of the event sequence from the burst tree, one can devise a model generating an event sequence with a hierarchical structure of bursts; for this, we first generate an ordinal burst tree and then the event sequence by incorporating a set of IETs drawn from the IET distribution.
@@ III.1 Kernel-based modeling
Let us begin with $n$ events or leaf nodes, each of which makes a burst of size one in the beginning. Their timings are to be calculated after the ordinal burst tree derived from the kernel is associated with the IET distribution. To describe the merging process in parallel with the coagulation process [33, 34, 35, 36, 37, 38, 39], we introduce an auxiliary time step $s$ that simply counts the number of merges in the merging process. We denote the burst size distribution at the time step $s$ by $Q_{s}(b)$. Then, at the initial time step $s=0$, one gets $Q_{0}(b)=\delta_{b,1}$, where $\delta$ is Kronecker delta. The merging process is as follows:
@@
We consider five burst-merging kernels. The first three are constant, additive, and product kernels as follows:
@@
Both $K_{\rm add}$ and $K_{\rm prod}$ are increasing functions of $b$ and $b^{\prime}$, implying that the larger burst is more likely to be chosen for the merge. Therefore, these kernels implement the preferential mixing of bursts that is responsible for heavy-tailed burst size distributions. In contrast, since all the kernels introduced do not require any similarity of bursts to be merged, they may not lead to the assortative mixing of bursts. Here the assortative mixing implies a tendency that bigger (smaller) bursts follow bigger (smaller) bursts [40], accounting for positive correlations between consecutive burst sizes. As the empirical results indicate both preferential and assortative mixing of bursts [20], we employ two more kernels dictating both preferential and assortative mixing:
$K_{\rm emp1}(b,b^{\prime})=$ $\frac{b+b^{\prime}}{1+c_{1}(b-b^{\prime})^{2}},$ (4)
$K_{\text{emp2}}(b,b^{\prime})=$ $\left[1+c_{2}(\ln b+\ln b^{\prime})\right]$
×$\times\left[1+c_{3}e^{-(\ln b-\ln b^{\prime})^{2}/c_{4}}\right],$ (5)
where $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$ are positive parameters. Here $K_{\rm emp1}$’s numerator and denominator are for preferential and assortative mixing, respectively. The first and second parentheses of $K_{\rm emp2}$ indicate preferential and assortative mixing, respectively. Note that $K_{\rm emp2}$ has been proposed in Ref. [20]. Finally, we remark that all five kernels considered are symmetric with respect to their arguments, namely, $K(b,b^{\prime})=K(b^{\prime},b)$, implying the time-reversal symmetry of the generated ordinal burst tree.
Once the ordinal burst tree $\mathcal{G}$ is constructed, we then randomly draw $n-1$ IETs from $P(\tau)$ to get a set of $n-1$ IETs. We adopt a power-law IET distribution defined over the values of $\tau\in\{1,2,\ldots,\tau_{\rm c}\}$ as follows:
@@
where $\tau_{\rm c}\gg 1$ is the upper bound of the IET. We assign each of the $n-1$ IETs, in a descending order from the largest to the smallest IETs, to the internal nodes, labeled from rank $1$ to rank $n-1$, respectively. Note that the root node with rank one has the largest IET $\tau_{\text{max}}$. After setting $t_{0}=1$, we iteratively calculate the event timings using the recursive formula $t_{i}=t_{i-1}+\hat{\tau}_{u(i)}$, where $u(i)$ denotes the $i$th visited internal node when traversing the ordinal burst tree in inorder.
@@ III.2 Measures for higher-order temporal correlations
Here $\mu_{1}$ ($\mu_{2}$) and $\sigma_{1}$ ($\sigma_{2}$) are the mean and standard deviation of $B_{\Delta t}$ without the last (first) burst, respectively. $M_{\Delta t}$ has a value in the range of $[-1,1]$; $M_{\Delta t}$ may have a positive (negative) value if consecutive burst sizes are positively (negatively) correlated with each other. If the burst sizes are uncorrelated, $M_{\Delta t}$ is close to zero.
@@
We remark that thanks to the analogy between the merging process and the coagulation process, exact solutions of $Q_{s}(b)$ are available for the constant, additive, and product kernels [39]. In the case of the constant kernel, the cluster size distribution in the coagulation process translates into the burst size distribution as follows:
@@
where $b_{s}\equiv\frac{t(s)}{2}$. Here the time variable $t$ in the coagulation process is an increasing function of $s$ in the merging process but its functional form is non-trivial. Since $b_{s}$ in Eq. (11) is the same as the average burst size, which is $\frac{n}{n-s}$, one gets $t(s)=\frac{2n}{n-s}$. The exponential burst size distribution implies that there is no correlation between IETs [22]. In the case of the additive kernel, we can write
@@
where $b^{\prime}_{s}\equiv 2e^{t(s)}$. Again, $t$ in the coagulation process is a function of $s$ in the merging process. The burst size distribution shows a power-law regime with exponent $\beta=3/2$. Finally, in the case of the product kernel, we get
@@
where $b^{\prime\prime}_{s}\equiv\frac{1}{t(s)-1-\log t(s)}$. In the context of the coagulation process, a gelation transition occurs at $t(s)=1$, when a particle of infinite size or a gel appears [39]. This implies in the merging process that a very big burst emerges at a finite time step $s$, satisfying $t(s)=1$. The burst size distribution shows a power-law regime with exponent $\beta=5/2$ for $t(s)<1$, while it has a power-law tail with the same exponent for $t(s)\geq 1$. These analytical results will be discussed against the simulation results in the next Section.
Figure 2: Visualization of event sequences with $n=300$ events generated using the constant, additive, and product kernels as well as two kernels inspired by empirical results with $c_{1}=0.02$, $c_{2}=3$, $c_{3}=100$, and $c_{4}=4$ in Eqs. (1)–(5) (from top to bottom). Here we have used exactly the same set of 299 IETs drawn from $P(\tau)\propto\tau^{-3.5}$ for every event sequence.
@@
Figure 3: Simulation results of the burst size distributions $Q_{s}(b)$ and $Q_{\Delta t}(b)$, memory coefficients for bursts $M_{\Delta t}$ in Eq. (8), and autocorrelation functions $A(t_{\rm d})$ in Eq. (9) (from left to right) obtained from $50$ event sequences of $n=5\cdot 10^{5}$ events generated using $K_{\rm const}$, $K_{\rm add}$, $K_{\rm prod}$, $K_{\rm emp1}$ with $c_{1}=0.02$, and $K_{\rm emp2}$ with $c_{2}=3$, $c_{3}=100$, and $c_{4}=4$ [Eqs. (1)–(5)] (from top to bottom). $\langle b\rangle_{s}$ and $\langle b\rangle_{\Delta t}$ are average burst sizes of $Q_{s}(b)$ and $Q_{\Delta t}(b)$, respectively. As for $Q_{\Delta t}(b)$, we show results for $\alpha=2.5$. Black lines with exponent values in panels (e) and (i) are for guiding eyes. Black lines in panels (n) and (r) show fitting ranges and results of the burst size exponent $\beta$ for the curves with $\Delta t=64$. Black lines in the rightmost panels show fitting ranges and results of the decay exponent $\gamma$ in Eq. (10), summarized in Table 1. Error bars in $M_{\Delta t}$ and $A(t_{\rm d})$ denote standard errors.
IV Results

### v2

@@ Abstract
Understanding characteristics of temporal correlations in time series is crucial for developing accurate models in natural and social sciences. The burst-tree decomposition method was recently introduced to reveal higher-order temporal correlations in time series in a form of an event sequence, in particular, the hierarchical structure of bursty trains of events for the entire range of timescales [Jo et al., Sci. Rep. 10, 12202 (2020)]. Such structure has been found to be simply characterized by the burst-merging kernel governing which bursts are merged together as the timescale for defining bursts increases. In this work, we study the effects of kernels on the higher-order temporal correlations in terms of burst size distributions, memory coefficients for bursts, and the autocorrelation function. We employ several kernels, including the constant, sum, product, and diagonal kernels as well as those inspired by empirical results. We generically find that kernels with preferential merging lead to heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the interevent time distribution. In addition, thanks to the analogy to the coagulation process, analytical solutions of burst size distributions for some kernels could be obtained. Our findings may shed light on the role of burst-merging kernels as underlying mechanisms of higher-order temporal correlations in time series.
@@ I Introduction
Recently, Jo et al. introduced a novel method called burst-tree decomposition to analyze temporal correlations in the event sequence [20]. This method represents an event sequence in terms of the burst tree, which is decomposed into a set of IETs and an ordinal burst tree. The ordinal burst tree exactly captures the hierarchical structure of correlations between IETs in a systematic way. It turns out that several time series data from diverse backgrounds with different IET distributions show the similar burst-tree structure in terms of heavy-tailed burst size distributions and positive correlations between consecutive burst sizes for a wide range of timescales. The latter was quantified in terms of a memory coefficient for bursts. Originally, the memory coefficient has been proposed for detecting correlations between two consecutive IETs [17]. The burst-tree structure could be simply summarized by a burst-merging kernel governing which bursts are merged together as the timescale increases. The estimated burst-merging kernels from the data have revealed the preferential and assortative mixing structure of bursts. Then a model adopting a simple mathematical form of the burst-merging kernel based on the empirical analysis was studied to successfully generate temporal correlations in the original time series, providing a powerful tool for modeling and analyzing complex systems.
@@
The modeling approach using burst-merging kernels is analogous to the coagulation process by which unit particles are merged according to some selection rules or kernels, to become clusters [33, 34, 35, 36, 37, 38, 39, 40, 41]. Different approaches have been taken to derive cluster size distributions as well as power-law exponent values in the cases of power-law distributions of cluster sizes. In particular, Wattis employed a discrete deterministic mean-field approach to derive the analytical solution of the cluster size distribution for three solvable kernels, namely, constant, sum, and product kernels [41]. Similar results for the diagonal kernel have been obtained in Ref. [40]. Since unit particles and clusters in the coagulation process respectively correspond to individual events and bursts in the event sequence, the analytical solutions of cluster size distributions in the coagulation process can be translated into the burst size distributions of the event sequence generated using the same kernels. It implies that the analytical solutions of cluster size distributions in Refs. [41] and [40] can be used to confirm the burst size distributions obtained when the constant, sum, product, and diagonal kernels are employed as the burst-merging kernels.
@@
Despite the importance of burst-merging kernels in time series analysis, their impact on higher-order temporal correlations remains poorly understood. To address this issue, we conduct a systematic comparison of various kernels, including the constant, sum, product, and diagonal kernels as well as those inspired by empirical results. By analyzing the event sequences generated using these kernels and power-law IET distributions as inputs, we investigate burst size distributions, memory coefficients for bursts, and ACFs. We generically find that kernels with preferential merging lead to the heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the IET distribution. These findings provide new insights into the role of burst-merging kernels as underlying mechanisms of higher-order temporal correlations in time series.
@@ II Burst-tree decomposition method
The burst-tree decomposition method was recently introduced to systematically reveal temporal correlations in time series, in particular, the hierarchical structure of bursts for a wide range of timescales defining the bursts [20]. A burst refers to a sequence of events that occur in rapid succession, followed by a period of relative inactivity [8]. To be precise, for a given timescale $\Delta t$, a burst is defined as a set of events where interevent times (IETs) between consecutive events in the burst are less than or equal to $\Delta t$, whereas the IETs between bursts are greater than $\Delta t$ [22]. It implies that two consecutive bursts detected using a timescale $\Delta t$ would be merged for a bigger timescale $\Delta t^{\prime}$ ($>\Delta t$) if $\Delta t^{\prime}$ is larger than the IET separating these two bursts. Such merging pattern is exactly captured by the burst-tree decomposition method.
@@
Since the burst tree contains exactly the same information in the event sequence, one can recover the original event sequence $\mathcal{E}$ from the derived burst tree $\mathcal{T}$ too. This is done by performing an inorder traversal of the burst tree. If the $i$th visited internal node is indexed by $u$, we denote it by $u(i)$. Then the event timings are recursively computed by $t_{i}=t_{i-1}+\hat{\tau}_{u(i)}$. Here $t_{0}$ can be taken from the time series data, or assumed to be 00 for simplicity. Thanks to the derivability of the event sequence from the burst tree, one can devise a model generating an event sequence with a hierarchical structure of bursts; for this, we first generate an ordinal burst tree and then the event sequence by incorporating a set of IETs drawn from the IET distribution.
@@ III.1 Kernel-based modeling
Let us begin with $n$ events or leaf nodes, each of which makes a burst of size one in the beginning. Their timings are to be calculated after the ordinal burst tree derived from the kernel is associated with the IET distribution. To describe the merging process in parallel with the coagulation process [33, 34, 35, 36, 37, 38, 39, 40, 41], we introduce an auxiliary time step $s$ that simply counts the number of merges in the merging process. We denote the burst size distribution at the time step $s$ by $Q_{s}(b)$. Then, at the initial time step $s=0$, one gets $Q_{0}(b)=\delta_{b,1}$, where $\delta$ is Kronecker delta. The merging process is as follows:
@@
We consider seven burst-merging kernels. The first three are constant, sum, and product kernels defined as follows:
@@
Both $K_{\rm sum}$ and $K_{\rm prod}$ are increasing functions of $b$ and $b^{\prime}$, implying that larger bursts are more likely to be chosen for the merge than smaller bursts. Therefore, these kernels indicate the preferential merging of bursts that is believed to be responsible for heavy-tailed burst size distributions. In contrast, since all three kernels do not require any similarity of bursts to be merged, they may not lead to the assortative pattern of bursts. Here the assortativity implies a tendency that bigger (smaller) bursts follow bigger (smaller) bursts [42], accounting for positive correlations between consecutive burst sizes. To implement assortative merging of bursts, we consider a diagonal kernel as
$K_{\rm diag}(b,b^{\prime})=\delta_{b,b^{\prime}}.$ (4)
With this kernel, only bursts of the same size can merge to become a bigger burst. Thus, one can expect the burst sizes to be in the form of $2^{k}$ for non-negative integers $k$. In addition, for the burst containing all events to exist, $n$ must be of the form of $2^{k}$.
As the empirical results indicate both preferential and assortative merging of bursts [20], we employ three more kernels dictating both preferential and assortative merging behaviors:
$K_{{\rm diag\text{-}pre}}(b,b^{\prime})=b\,\delta_{b,b^{\prime}},$ (5)
$K_{{\rm sum\text{-}diag}}(b,b^{\prime})=\frac{b+b^{\prime}}{1+c_{1}(b-b^{\prime})^{2}},$ (6)
$K_{\rm emp}(b,b^{\prime})=\left[1+c_{2}\ln(bb^{\prime})\right]\left[1+c_{3}e^{-c_{4}(\ln b-\ln b^{\prime})^{2}}\right],$ (7)
where $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$ are non-negative parameters. The diagonal-preferential kernel $K_{{\rm diag\text{-}pre}}$ is the diagonal kernel $K_{\rm diag}$ in Eq. (4) multiplied by $b$, hence it is an increasing function of $b$ or $b^{\prime}$ along the diagonal line $b=b^{\prime}$. Next, setting $c_{1}=0$ in $K_{{\rm sum\text{-}diag}}$ reduces it to $K_{\rm sum}$; the denominator of $K_{{\rm sum\text{-}diag}}$ with positive $c_{1}$ implements the assortative merging of bursts. Finally, the first and second parentheses of $K_{\rm emp}$ indicate preferential and assortative merging, respectively. The functional complexity of $K_{\rm emp}$ is indeed inherited from the empirically observed kernel [20], hence we have suggested $K_{{\rm sum\text{-}diag}}$ as a simpler version of $K_{\rm emp}$. We also remark that all the kernels considered are symmetric with respect to their arguments, namely, $K(b,b^{\prime})=K(b^{\prime},b)$, implying the time-reversal symmetry of the generated ordinal burst tree.
Once the ordinal burst tree $\mathcal{G}$ is constructed, we then randomly draw $n-1$ IETs from $P(\tau)$ to get a set of $n-1$ IETs. We adopt a power-law IET distribution with the exponent $\alpha$, defined over the values of $\tau\in\{1,2,\ldots,\tau_{\rm c}\}$, as follows:
@@
where $\tau_{\rm c}\gg 1$ is the upper bound of the IET. We assign each of the $n-1$ IETs, in a descending order from the largest to the smallest IETs, to the internal nodes, labeled from rank $1$ to rank $n-1$, respectively. Note that the root node with rank $1$ has the largest IET $\tau_{\text{max}}$. After setting $t_{0}=1$, we iteratively calculate the event timings using the recursive formula $t_{i}=t_{i-1}+\hat{\tau}_{u(i)}$, where $u(i)$ denotes the $i$th visited internal node when traversing the ordinal burst tree in inorder.
@@ III.2 Measures for higher-order temporal correlations
Here $\mu_{1}$ ($\mu_{2}$) and $\sigma_{1}$ ($\sigma_{2}$) are the average and standard deviation of $B_{\Delta t}$ without the last (first) burst, respectively. $M_{\Delta t}$ has a value in the range of $[-1,1]$; $M_{\Delta t}$ may have a positive (negative) value if consecutive burst sizes are positively (negatively) correlated with each other. If the burst sizes are uncorrelated, $M_{\Delta t}$ is close to zero.
Figure 2: Visualization of event sequences with $n=256$ events generated using the constant, sum, product, and diagonal kernels as well as kernels inspired by empirical results with $c_{1}=0.01$, $c_{2}=3$, $c_{3}=100$, and $c_{4}=0.25$ in Eqs. (1)–(7) (from top to bottom). Each vertical line depicts an event. We have used exactly the same set of 255 interevent times drawn from $P(\tau)\propto\tau^{-3.5}$ for all cases.
@@
III.3 Analytical results of burst size distributions
Thanks to the analogy between the merging process and the coagulation process, exact solutions of $Q_{s}(b)$ are available for the constant, sum, and product kernels [41]. In the case of the constant kernel, the cluster size distribution in the coagulation process translates into the burst size distribution as follows:
@@ III.3 Analytical results of burst size distributions
where $b_{s}\equiv\frac{t(s)}{2}$. Here the time variable $t$ in the coagulation process is an increasing function of $s$ in the merging process but its functional form is not always trivial to obtain. Since $b_{s}$ in Eq. (13) is the same as the average burst size, which is $\frac{n}{n-s}$, one gets $t(s)=\frac{2n}{n-s}$. The exponential burst size distribution implies that there is no correlation between IETs [22]. In the case of the sum kernel, we can write
@@
where $b^{\prime}_{s}\equiv 2e^{t(s)}$. Again, $t$ in the coagulation process is a function of $s$ in the merging process. The burst size distribution shows a power-law regime with exponent $\beta=3/2$. Next, in the case of the product kernel, we get
@@
where $b^{\prime\prime}_{s}\equiv\frac{1}{t(s)-1-\log t(s)}$. In the context of the coagulation process, a gelation transition occurs at $t(s)=1$, when a particle of infinite size or a gel appears [41]. This implies in the merging process that a very big burst emerges at a finite time step $s$, satisfying $t(s)=1$. The burst size distribution shows a power-law regime with exponent $\beta=5/2$ for $t(s)<1$, while it has a power-law tail with the same exponent for $t(s)\geq 1$.
Finally, the diagonal-preferential kernel $K_{{\rm diag\text{-}pre}}$ can be considered as a special case of the form of $K(x,y)=x^{1+\lambda}\delta(x-y)$ that was studied in Ref. [40], where $\lambda\geq 0$, and $x$ and $y$ are continuous variables. The analytical solution of the cluster size distribution for the case with $\lambda=0$ in Ref. [40] can be translated into the burst size distribution as
$Q_{s}(b)\propto\frac{b^{-1}}{t(s)}\left[1-c\left(\frac{b}{t(s)}\right)^{\Delta}\right],$ (16)
where $c>0$ is a constant determined by the normalization, and $\Delta\approx 2.69$. These analytical results will be compared to the simulation results in the next Section, but indirectly via $Q_{\Delta t}(b)$. It is because once the number of bursts for a given $\Delta t$ is determined as $m_{\Delta t}$, the corresponding auxiliary time step $s$ is simply given as $n-m_{\Delta t}$. Then, since $Q_{s=n-m_{\Delta t}}(b)=Q_{\Delta t}(b)$, it is enough to show the simulation results of $Q_{\Delta t}(b)$ to see if the analytic solutions predict the simulations well.
@@
Figure 3: Simulation results of burst size distributions $Q_{\Delta t}(b)$, memory coefficients for bursts $M_{\Delta t}$ in Eq. (10), and autocorrelation functions $A(t_{\rm d})$ in Eq. (11) (from left to right) obtained from $50$ event sequences of $n=5\cdot 10^{5}$ events generated using $K_{\rm const}$, $K_{\rm sum}$, and $K_{\rm prod}$ [Eqs. (1)–(3)] (from top to bottom). In panels (a, d, g), we show $Q_{\Delta t}(b)$ for several values of $\Delta t$ in the case with the IET exponent $\alpha=2.5$, and $\langle b\rangle_{\Delta t}$ is the average burst size of $Q_{\Delta t}(b)$. Black lines with exponent values in panels (d, g) are for guiding eyes. Black lines in panels (c, f, i) show fitting ranges and results of the decay exponent $\gamma$ in Eq. (12), summarized in Table 1. Error bars in $M_{\Delta t}$ and $A(t_{\rm d})$ denote standard errors.
IV Numerical results